OpenStudy (anonymous):

hello there i have some questions about logarithm functions here its log a-2 logb+3logc

hartnn (hartnn):

you want to simplify this ? \(\large \log a-2 \log b+3\log c\)

hartnn (hartnn):

here are some log rules you will need \(\large \log x +\log y = \log xy \\ \large \log x - \log y = \log (x/y ) \\ \large x\log y = \log y^x \)

hartnn (hartnn):

using the last one, we get \(\large 2 \log b = \log b^2 \\ \large 3 \log c = \log c^3 \)

hartnn (hartnn):

and using the first two, we get \(\huge \log (\dfrac{ac^3}{b^2})\) thats it!